how much 10% acid solution should be added to 30 liters of 25% acid solution to achieve a 15% solution?
Answer by stanbon(75887)

how much 10% acid solution should be added to 30 liters of 25% acid solution to achieve a 15% solution?
------
acid + acid = acid
0.10x + 0.25*30 = 0.15(x+30)
--------
10x + 25*30 = 15x + 15*30
5x = 10*30
x = 60 liters (amt. of 10% solution needed)
============================================
Cheers,
Stan H.
=================